OpenStudy (anonymous):

PLEASE HELP WITH THIS VECTORS PROBLEM! Determine each resultant: a) 10 km/h north followed by 8 km/h west. b) 25 N at 10 degrees to the horizontal and 30 N horizontally Please show all work, Thanks (:

OpenStudy (anonymous):

do you also need the direction from north?

OpenStudy (anonymous):

I dont know, this is the question :P

OpenStudy (anonymous):

Better safe then sorry

OpenStudy (anonymous):

 can you use the pythagorean theorem here?

OpenStudy (anonymous):

\[8^{2}+10^{2} =\sqrt164\] \[=12.81\]

OpenStudy (anonymous):

yep... or you can leave it as \(\large \sqrt{164}=2\sqrt{41} \). now the direction.... notice \(\large tan\theta=\frac{8}{10}\) can you solve for theta?

OpenStudy (anonymous):

\[\Theta=\tan ^{-1}(0.8)\]?

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

i got 38.66 degrees west of north should be the direction.

OpenStudy (anonymous):

Same here